# How faithful are RAG models? Quantifying the tug-of-war between RAG and LLMs’ internal prior

### v1

@@ 3.1 Concordance
In Figure 2 (left side plots), we observe a consistent negative relationship between the token probability of the model’s prior answer and the associated RAG preference rate for all six QA datasets. To visualize an even distribution across probabilities, we bin the probabilities into ten equidistant bins in the range of [0.0,1.0]0.01.0[0.0,1.0][ 0.0 , 1.0 ]. We additionally present the slope from performing a linear regression on the binned probability values against the RAG preference rate in Table 1. The slope indicates the effect of stronger model confidence on the model’s preference for the information presented in the retrieved context; we observe different slopes (ranging from -0.1 to -0.45), suggesting that the effectiveness of RAG in different QA domains can be characterized as being relatively susceptible (e.g., with Dates questions) or robust (e.g., with News questions) to the model’s internal prior knowledge confidence. Specifically, a slope of -0.45, for instance, can be interpreted as expecting a 4.5% decrease in the likelihood of the LLM preferring the contextual information for every 10% increase in the probability of the model’s prior response.
3.2.1 RAG Preference Rate vs Deviation from Prior
We also consider the degree of deviation between the model’s prior response and the value contained in the retrieved context (Figure 2, right side plots). A similar pattern emerges in this analysis: as the RAG value diverges from the model’s prior, the model is less likely to adopt the RAG value over its own initial response. We additionally plot data split into the upper and lower half percentiles and observe that, across all six datasets, the lower probability prior responses are monotonically lower than the higher probability response tokens. Thus, the correlation between deviation and RAG response rate holds across both bands of probabilities.
3.2.2 Effect of prompting technique on RAG adherence
To assess the degree of influence that the specific prompting technique has on RAG adherence, we test two additional prompts (”strict” and ”loose”) on GPT-4. The strict prompt is intended to coerce the model to disregard its own prior response, while the loose prompt is intended for the model to arbitrate between its own prior and the contextual information provided. In Figure 4, the strict prompt has uniformly higher RAG adherence than the standard prompt. The loose prompt, on the other hand, results in much lower RAG adherence rates as prior probability increases. Interestingly, the slope is also steeper, indicating a larger per-unit decrease in RAG preference as the prior probability increases. The choice of prompt is thus an important mechanism for influencing the LLM’s RAG preferences.
3.2.3 Differences in effects between GPT-4, GPT-3.5, and Mistral-7B
We report the same analyses when using GPT-3.5 and Mistral-7B in Table 2 and Figure 5. We observe significantly lower performance both in concordance of the prior and with RAG. However, as seen in Figure 5, we nonetheless observe the same inverse trends in these two models as seen in the results with GPT-4. Of note, some datasets (like Latest News) perform poorly without RAG (the model refused the vast majority of queries or provided invalid responses), and thus the prior token probabilities could not be analyzed. In the Mistral-7B results, we also observe that some of the responses using RAG could not consistently provide valid responses.
GPT-4 Concordance (Prior) Concordance (w/ RAG) Slope
Drug Dosage 0.554 0.884 -0.26
Sports Stats 0.240 0.943 -0.18
Latest News 0.133 0.936 -0.10
Wikipedia Dates 0.433 0.995 -0.45
Wikipedia Names 0.350 0.965 -0.13
Wikipedia Locations 0.375 0.920 -0.28
Average 0.347 0.940 -0.23
Table 1: Concordance between the GPT response and the reference values for each dataset. Prior refers to GPT-4 responses without context, and ”w/ RAG” refers to responses with the relevant retrieved context included in the prompt. Additionally, we include the slope of the relationship between prior probability and RAG preference rate. For instance, the average slope is -0.23, which means that for every 10% increase in the probability of the prior token, we observe a 2.3% decreased likelihood of RAG preference.
@@ 4 Discussion
While RAG is becoming standard practice in commercially available LLMs, the reliability of such systems is still understudied. Our experiments uncover several mechanisms that modulate the degree to which LLMs adhere to RAG systems. Specifically, we quantify a tug-of-war between the strength of the model’s prior and the rate at which the model adheres to the RAG document’s facts. This effect is at odds with claims that RAG itself can fix hallucinations alone, and occurs even when the model is prompted to adhere to RAG documents strictly.
RAG systems have a unique appeal over traditional search engines in that they can incorporate prior knowledge to fill in the gaps and extrapolate the retrieved information. We find that this comes with trade-offs – namely, that such priors can override information provided in documents. When perturbing RAG documents over a wide interval of values, the point at which models revert to their prior responses, or ”tipping points”, are latent and heterogeneous across different models and domains. While strong priors are not inherently problematic (and can often serve to safeguard models), the lack of explicit expectations around how models will mix reference documents with their priors can lead to downstream issues. For example, if RAG systems are used to extract nested financial data to be used in an algorithm, what will happen if there is a typo in the financial documents? Will the model notice the error and if so, what data will it provide in its place? Given that LLMs are soon to be widely deployed in many domains including medicine and law, users and developers alike should be cognizant of their unintended effects, especially if users have preconceptions that RAG-enabled systems are, by nature, always truthful.
There are several key limitations in our analyses. First, RAG systems can be deployed to many more domains than can be covered by our analyses. However, we hope that our study ixix domains paints an initial picture of the nature of RAG systems. Second, to make our experiments tractable, our question-generation process is strictly fact-based and does not require multi-step logic, document synthesis, or other higher-level reasoning. Third, the perturbations we produce are based on our priors for what would constitute a reasonable or unreasonable range of values. In a natural setting, we would imagine more discrete types of errors (eg. typos, ambiguities, missing information, etc.) which are harder to simulate. We also perform evaluations on GPT-3.5 and GPT-4 because the OpenAI API allows for access to token-wise log probabilities along with the responses. As a consequence, we are limited against performing more comprehensive evaluations on models such as Gemini and Claude because the APIs for these models do not provide access to such information.
LLMs are now commonly used as parts of larger, more complex systems. It is crucial to understand how these models interact with information with varying degrees of trustworthiness, accuracy, and uniformity. Our analysis shows that further work is required to characterize the risks of using LLMs to answer questions given contextual information. In particular, we find that model behavior can be erratic and unpredictable when presented with information that exists at the margin of its prior beliefs.
@@ Appendix A Appendix
Figure 5: Analyses for RAG preference rate against prior probability and deviation, using GPT-4 (blue), GPT-3.5 (orange), and Mistral-7B (green). Please see Figure 2 for full figure descriptions. Of note, some models did not generate any meaningful prior responses (due to refusal, improper responses, etc.) for certain datasets and thus could not be analyzed.
GPT-3.5 Concordance (Prior) Concordance (w/ RAG) Slope
Drug Dosage 0.052 0.509 -0.20
Sports Stats 0.005 0.599 -0.09
Latest News 0.008 0.839 N/A
Wikipedia Dates 0.275 0.985 -0.71
Wikipedia Names 0.285 0.965 -0.11
Wikipedia Locations 0.410 0.930 -0.16
Average 0.173 0.805 -0.25
Mistral-7B Concordance (Prior) Concordance (w/ RAG) Slope
Drug Dosage 0.550 0.677 -0.82
Sports Stats 0.240 0.057 N/A
Latest News N/A N/A N/A
Wikipedia Dates 0.080 0.840 -0.77
Wikipedia Names 0.065 0.760 -0.14
Wikipedia Locations 0.490 0.690 -0.030
Average 0.285 0.605 -0.44
Table 2: Concordance and Slope with GPT-3.5 and Mistral-7B. Please see Table 1 for a full table description of analyses performed using GPT-4.
A.1 Prompts
Due to the length of the prompts, we have stored them in a CSV within open-access and anonymized repository: https://anonymous.4open.science/r/rag-tug-of-war-1E48/prompts.csv

### v2

@@ 4 Results
In Figure 4, we observe a consistent negative relationship between the token probability of the model’s prior answer and the associated RAG preference rate for all six QA datasets. To visualize an even distribution across probabilities, we bin the probabilities into ten equidistant bins in the range of $[0.0,1.0]$. The slope indicates the effect of stronger model confidence on the model’s preference for the information presented in the retrieved context; we observe different slopes (ranging from -0.1 to -0.45), suggesting that the effectiveness of RAG in different QA domains can be characterized as being relatively susceptible (e.g., with Dates questions) or robust (e.g., with News questions) to the model’s internal prior knowledge confidence. Specifically, a slope of -0.45, for instance, can be interpreted as expecting a 4.5% decrease in the likelihood of the LLM preferring the contextual information for every 10% increase in the probability of the model’s prior response.
Figure 4: We additionally observe an inverse relationship between the context preference rate (y-axis) and the model’s prior response probability (x-axis). Context preference rate is defined as the proportion of responses that align with the information presented in the prompt as context. The model’s prior response probability is computed from the average log probability of the response tokens queried without context. Each plot visualizes the prior probability (grouped into 10 bins) against the context preference rate, along with the best-fit trend line and slope. Models that allow access to token probabilities are shown.
4.3.1 Initial Methods for Improving Prior vs. Context Conflict Resolution
Based on our observations from the relationship between the token probabilities and the rates of preference for context, we posit that comparing token probabilities between $r(q)$ and $r(q|c)$ can improve the abilities of models to resolve conflicts. In Table 3, Token Probability Correction is done by comparing the mean token probabilities of the model’s response with and without context. If the probability is higher for the prior than the contextual response, then we use the model’s generation without context as its final response. Otherwise, we just use the response with context. We find that this method improves the overall accuracy of all three models with a moderate increase in the prior bias of each model. Next, we observe that the probability distributions between prior responses and context-given responses are uncalibrated, where context-given response probabilities are extremely right-tailed while prior probabilities are nearly uniform. As a simple adjustment, we compare the percentiles rather than raw probability scores of each score, or the Calibrated Token Probability Correction. We find that calibrated token probability correction improves all models’ overall accuracy by 14% and context bias by 20%. At the same time, this introduces more prior bias, from 2% to 8.5%. However, this method outperforms a baseline of randomly replacing the final response with its prior – at the same bias rate of 8.5%, the random baseline has an accuracy of 57.5% as compared to the 75.4% from the method. While this paper focuses on developing the ClashEval benchmark, these results suggest that probability calibration is a promising approach to reduce prior and context bias deserving further investigation. It also is a natural baseline for future methods.
Model Correction Accuracy $\uparrow$ Context Bias $\downarrow$ Prior Bias $\downarrow$
GPT-4o No correction (Baseline) 0.615 (0.595, 0.636) 0.304 (0.287, 0.321) 0.021 (0.014, 0.028)
Token Probability Correction 0.693 (0.672, 0.714) 0.194 (0.177, 0.210) 0.043 (0.032, 0.053)
Calibrated Token Prob. Correction 0.754 (0.733, 0.775) 0.107 (0.093, 0.122) 0.085 (0.072, 0.098)
GPT-3.5 No correction (Baseline) 0.539 (0.521, 0.557) 0.313 (0.298, 0.328) 0.028 (0.021, 0.036)
Token Probability Correction 0.596 (0.575, 0.616) 0.253 (0.237, 0.269) 0.056 (0.046, 0.067)
Calibrated Token Prob. Correction 0.701 (0.678, 0.722) 0.110 (0.098, 0.124) 0.147 (0.132, 0.164)
Llama-3 No correction (Baseline) 0.500 (0.483, 0.515) 0.264 (0.250, 0.279) 0.021 (0.015, 0.027)
Token Probability Correction 0.556 (0.537, 0.574) 0.235 (0.220, 0.249) 0.046 (0.037, 0.055)
Calibrated Token Prob. Correction 0.649 (0.627, 0.669) 0.111 (0.099, 0.122) 0.188 (0.173, 0.204)
Table 3: For models which provide token probabilities, we evaluate the accuracy, context bias, and prior bias under three conditions: (1) No correction, which is the baseline result from this paper, (2) the token probability correction, and (3) the calibrated token probability correction.
@@ 5 Discussion
The ClashEval benchmark dataset and evaluations provide novel insights into how LLMs arbitrate between their own internal knowledge and contextual information when the two are in conflict.
A key finding is that even the most advanced LLMs like GPT-4o exhibit a strong context bias, overriding their own correct prior knowledge over 60% of the time when presented with incorrect information in the retrieved documents. However, this bias is not absolute - the degree to which the retrieved content deviates from truth negatively correlates with the context preference rate. Interestingly, each LLM exhibits a different prior distribution over truthfulness across domains, such that the same perturbation level affects each model differently. For instance, for a given magnitude of deviation, Claude Opus adheres to incorrect contextual information 30% less often than GPT-4o. While GPT-4o achieves state-of-the-art results on general-purpose tasks, it exhibits higher context bias compared to smaller models like Claude Sonnet. This finding suggests that performance on knowledge-based benchmarks may not automatically mean it is most suitable for RAG settings. Additionally, we find that LLMs are calibrated to selectively defer to external evidence when they are less certain about a given query. However, each model differs in how well-calibrated they are. While strong priors are not inherently problematic, the lack of explicit expectations around how models will decide to use contextual information remains a risk. We propose a simple method for improving models under ClashEval, and hope that future work can improve upon this baseline.
Our analyses have several key limitations. First, RAG systems can be deployed to many more domains than can be covered by our analyses. Second, to make our experiments tractable, our question-generation process is strictly fact-based and does not require multi-step logic, document synthesis, or other higher-level reasoning. Third, our dataset contains an enriched rate of contextual errors, so the reported metrics are not meant to represent bias rates in the wild. Fourth, our proposed token probability method only applies to models which provide probability outputs. Finally, even though this dataset is intended to improve an LLM’s ability to provide users with accurate information, bad actors could use such information to exploit the shortcomings of certain models described in this paper.
As retrieval-augmented AI systems become increasingly prevalent, we hope our dataset and insights spur further research into improving the robustness and calibration of such models. Resolving the tension between parametric priors and retrieved information is a crucial challenge on the path to safe and trustworthy language models.
@@ Appendix A Appendix
Model Context Bias $\downarrow$ Prior Bias $\downarrow$ Accuracy $\uparrow$
Claude Opus 0.157 (0.141, 0.174) 0.021 (0.014, 0.029) 0.743 (0.723, 0.763)
Claude Sonnet 0.201 (0.184, 0.215) 0.025 (0.018, 0.033) 0.658 (0.641, 0.678)
Gemini 1.5 0.245 (0.231, 0.260) 0.037 (0.029, 0.046) 0.624 (0.607, 0.641)
GPT-4o 0.304 (0.287, 0.321) 0.021 (0.013, 0.028) 0.615 (0.594, 0.633)
GPT-3.5 0.313 (0.298, 0.329) 0.028 (0.021, 0.036) 0.539 (0.522, 0.558)
Llama-3 0.264 (0.250, 0.280) 0.021 (0.015, 0.027) 0.500 (0.482, 0.518)
Table 4: We compare six top-performing models across three metrics. Context bias is when the model chooses the context answer when its prior was correct. Prior bias is when the model chooses its prior when the context answer is correct. Finally, accuracy is a straightforward measure of the fraction of times it can produce the correct answer. We find that Claude Opus performs the best across all metrics with a context bias rate of 0.157.
Figure 5: We plot the data from Table 4 – each model’s performance across three metrics in different colors, along with 95% confidence intervals.
Figure 6: Effect of different prompts using GPT-4 on context preference rate vs prior probability. The "Strict" prompt strongly enforces literal adherence to the retrieved context, while the "Loose" prompt encourages the model to make a reasonable judgment in light of the provided context. We observe lower and steeper drops in context preference with the loose vs strict prompts, suggesting that prompt wording plays a significant factor in controlling context preference. Full prompts are provided in our GitHub repository.